# 4D Palatini–Cartan Gravity in Hamiltonian Form

### v1

@@ 3.2. The Hamiltonian action
Moreover, the Hamiltonian analysis in [CS2019, CCS2020] shows that $\frac{1}{2}\underline{{\boldsymbol{e}}}^{2}\partial_{\mathrm{n}}{\underline{{\boldsymbol{\omega}}}}$ is indeed the analogue of the $p\dot{q}$ term in Hamiltonian mechanics, in the sense that, at fixed $t\in I$,
@@
(3.12) $\underline{{\boldsymbol{e}}}\,\underline{{\mathrm{d}}}_{\underline{{\boldsymbol{\omega}}}}\underline{{\boldsymbol{e}}}$ $=0,$
(3.13) $\epsilon_{\mathrm{n}}\left(\underline{{\boldsymbol{e}}}\,\underline{F}_{{\underline{{\boldsymbol{\omega}}}}}+\frac{\Lambda}{3!}\underline{{\boldsymbol{e}}}^{3}\right)$ $=0,$
(3.14) $\underline{{\boldsymbol{e}}}_{i}\left(\underline{{\boldsymbol{e}}}\,\underline{F}_{{\underline{{\boldsymbol{\omega}}}}}+\frac{\Lambda}{3!}\underline{{\boldsymbol{e}}}^{3}\right)$ $=0.$
The first, together with the structural constraint (3.6a), is the torsion-free condition. The second and the third are known has the Hamiltonian and the momentum constraints, respectively. It is shown in [CCS2020] that these three constraints are first class. Moreover, the first generates gauge transformations, the second diffeomorphisms transversal to $\Sigma$, and the third diffeomorphism along $\Sigma$.
@@ 3.3. Initial and final conditions
The $1$-form $\alpha^{\Sigma}$ in (3.11) can in particular be considered at $\Sigma\times\{t_{0}\}$ and $\Sigma\times\{t_{1}\}$ and corresponds to the boundary terms appearing when computing the variation of the Hamiltonian action $S_{\mathrm{Ham}}$.
@@
In order to fix instead $\underline{{\boldsymbol{e}}}$ at the boundary, we have to change the $1$-form $\alpha^{\Sigma}$ by an exact term $\delta f^{\Sigma}$. The problem is that ${\underline{{\boldsymbol{\omega}}}}$ is not a well-defined $1$-form, for it is a connection. We can however write ${\underline{{\boldsymbol{\omega}}}}={\underline{{\boldsymbol{\omega}}}}_{0}+{\underline{{\boldsymbol{\omega}}}}^{\prime}$, which we regard as a change of variables ${\boldsymbol{\omega}}\mapsto{\boldsymbol{\omega}}^{\prime}$, where ${\underline{{\boldsymbol{\omega}}}}_{0}$ is a reference connection. Since ${\underline{{\boldsymbol{\omega}}}}_{0}$ is fixed, we have $\delta{\underline{{\boldsymbol{\omega}}}}_{0}=0$ and therefore $\alpha^{\Sigma}=\int_{\Sigma}\frac{1}{2}\underline{{\boldsymbol{e}}}^{2}\delta{\underline{{\boldsymbol{\omega}}}}^{\prime}$. We can now add to $S_{\mathrm{Ham}}$ (at $t=t_{0}$ and $t=t_{1}$, with appropriate signs) the term [^14] This is the general way of modifying a theory buy a boundary term in order to be able to pass from Neumann to Dirichlet boundary conditions. For PC, see [Corichi_2016, equation (4.7)] and references therein.
@@
This modifies the action, which we now denote by $S_{\mathrm{Ham}}^{\prime}$, in such a way that, when taking variations, the boundary terms are presently given by the $1$-form
@@ 4. Boundaries
In particular, if we pick $e_{\mathrm{n}}=\epsilon_{\mathrm{n}}$ and $\underline{{\boldsymbol{e}}}$ to correspond to Schwarzschild space, then $\mathbf{M}$ is the ADM [ADM1959] mass. [^17] For an overview of the ADM mass in the PC formalism, see [Ashtekar2008, OS2020bis]. Note that our discussion parallels that of [P04, Section 4.3] and references therein.
@@ 5. Euclidean gravity
    The reference $(0,1)$-form $\epsilon_{\text{n}}$ is normalized by $(\epsilon_{\text{n}},\epsilon_{\text{n}})=1$.
@@ 6. Towards quantization
To be more precise, one first has to take care of the symmetries of the PC action. If we do this in the BV formalism [BV1], see [CSPCH], we can then extend the splitting discussed in Section 3 to all the BV fields. At this point we can view the integration of the auxiliary field $\underline{{\boldsymbol{v}}}$ as a particular application of a BV pushforward [CMR2015, Section 2]. In the companion paper [CCfuture], we perform these steps and show this way that the PC theory on a cylinder is equivalent to the AKSZ [AKSZ] theory associated to its boundary BFV data constructed in [CCS2020b, Section 4]. (In [CCfuture], we will actually allow more general nondegeneracy conditions than in Definition 3.2 and will also consider the higher-dimensional case.)
The second step towards quantization would be to compute the Hamilton–Jacobi action of theory presented in this paper, following [CMW22], and to use the corresponding BV–AKSZ formulation to start computing the perturbative functional integral, following [CMW23]. The renormalization issues should start appearing at this point.

### v2

@@ 3.2. The Hamiltonian action
Moreover, the Hamiltonian analysis in [CS19c, CCS21c] shows that $\frac{1}{2}\underline{{\boldsymbol{e}}}^{2}\partial_{\mathrm{n}}{\underline{{\boldsymbol{\omega}}}}$ is indeed the analogue of the $p\dot{q}$ term in Hamiltonian mechanics, in the sense that, at fixed $t\in I$,
@@
(3.12) $\underline{{\boldsymbol{e}}}\,\underline{{\mathrm{d}}}_{\underline{{\boldsymbol{\omega}}}}\underline{{\boldsymbol{e}}}$ =0=0,= 0 ,
(3.13) $\epsilon_{\mathrm{n}}\left(\underline{{\boldsymbol{e}}}\,\underline{F}_{{\underline{{\boldsymbol{\omega}}}}}+\frac{\Lambda}{3!}\underline{{\boldsymbol{e}}}^{3}\right)$ =0=0,= 0 ,
(3.14) $\underline{{\boldsymbol{e}}}_{i}\left(\underline{{\boldsymbol{e}}}\,\underline{F}_{{\underline{{\boldsymbol{\omega}}}}}+\frac{\Lambda}{3!}\underline{{\boldsymbol{e}}}^{3}\right)$ =0=0.= 0 .
The first, together with the structural constraint (3.6a), is the torsion-free condition. The second and the third are known has the Hamiltonian and the momentum constraints, respectively. It is shown in [CCS21c] that these three constraints are first class. Moreover, the first generates gauge transformations, the second diffeomorphisms transversal to $\Sigma$, and the third diffeomorphism along $\Sigma$.
@@ 3.3. Initial and final conditions
The 111-form $\alpha^{\Sigma}$ in (3.11) can in particular be considered at $\Sigma\times\{t_{0}\}$ and $\Sigma\times\{t_{1}\}$ and corresponds to the boundary terms appearing when computing the variation of the Hamiltonian action $S_{\mathrm{Ham}}$.
@@
In order to fix instead $\underline{{\boldsymbol{e}}}$ at the boundary, we have to change the 111-form $\alpha^{\Sigma}$ by an exact term $\delta f^{\Sigma}$. The problem is that ${\underline{{\boldsymbol{\omega}}}}$ is not a well-defined 111-form, for it is a connection. We can however write ${\underline{{\boldsymbol{\omega}}}}={\underline{{\boldsymbol{\omega}}}}_{0}+{\underline{{\boldsymbol{\omega}}}}^{\prime}$, which we regard as a change of variables ${\boldsymbol{\omega}}\mapsto{\boldsymbol{\omega}}^{\prime}$, where ${\underline{{\boldsymbol{\omega}}}}_{0}$ is a reference connection. Since ${\underline{{\boldsymbol{\omega}}}}_{0}$ is fixed, we have $\delta{\underline{{\boldsymbol{\omega}}}}_{0}=0$ and therefore $\alpha^{\Sigma}=\int_{\Sigma}\frac{1}{2}\underline{{\boldsymbol{e}}}^{2}\delta{\underline{{\boldsymbol{\omega}}}}^{\prime}$. We can now add to $S_{\mathrm{Ham}}$ (at $t=t_{0}$ and $t=t_{1}$, with appropriate signs) the term [^14] This is the general way of modifying a theory buy a boundary term in order to be able to pass from Neumann to Dirichlet boundary conditions. For PC, see [CRV16a, equation (4.7)] and references therein.
@@
This modifies the action, which we now denote by $S_{\mathrm{Ham}}^{\prime}$, in such a way that, when taking variations, the boundary terms are presently given by the 111-form
@@ 4. Boundaries
In particular, if we pick $e_{\mathrm{n}}=\epsilon_{\mathrm{n}}$ and $\underline{{\boldsymbol{e}}}$ to correspond to Schwarzschild space, then $\mathbf{M}$ is the ADM [ADM59a] mass. [^17] For an overview of the ADM mass in the PC formalism, see [AES08a, OS20a]. Note that our discussion parallels that of [Poi04a, Section 4.3] and references therein.
@@ 5. Euclidean gravity
    The reference (0,1)(0,1)( 0 , 1 )-form $\epsilon_{\text{n}}$ is normalized by $(\epsilon_{\text{n}},\epsilon_{\text{n}})=1$.
@@ 6. Towards quantization
To be more precise, one first has to take care of the symmetries of the PC action. If we do this in the BV formalism [BV77a], see [CSPCH], we can then extend the splitting discussed in Section 3 to all the BV fields. At this point we can view the integration of the auxiliary field $\underline{{\boldsymbol{v}}}$ as a particular application of a BV pushforward [CMR2015, Section 2]. In the companion paper [CC25a], we perform these steps and show this way that the PC theory on a cylinder is equivalent to the AKSZ [Ale+97a] theory associated to its boundary BFV data constructed in [CCS21b, Section 4]. (In [CC25a], we will actually allow more general nondegeneracy conditions than in Definition 3.2 and will also consider the higher-dimensional case.)
The second step towards quantization would be to compute the Hamilton–Jacobi action of theory presented in this paper, following [CMW22b], and to use the corresponding BV–AKSZ formulation to start computing the perturbative functional integral, following [CMW22c]. The renormalization issues should start appearing at this point.